# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 4 Methodology Perspective
4.1 Overview
The integration of deep learning techniques into urban computing has facilitated the development of various deep-learning-based data fusion methods. These techniques aim to leverage the inherent connections within diverse urban data streams. To better understand the differences in the underlying concepts of data fusion in these studies, we categorize them into four distinct categories, as illustrated in Figure 7. The definitions of each category are outlined as follows:
@@ 4.1 Overview
Multiplication (Sec.4.2.1) [DeepST Zhang et al. (2016), Bai et al. (2019b), ROD-Revenue Guo et al. (2019), leaf, text width=20em] ] [Feature Concatenation
(Sec.4.2.2) [ST-ResNet Zhang et al. (2017), DeepCrime Huang et al. (2018), DeepSTN+ Lin et al. (2019),
ST-MetaNet+ Pan et al. (2020), Liang et al. (2021), Chen et al. (2021b),
ST-SHN Xia et al. (2021), Huang et al. (2023a), Yin et al. (2023b), leaf, text width=20em] ] [Graph-based Feature
Fusion (Sec.4.2.3) [Multi-view Graph [Du et al. (2019); Geng et al. (2019a),
Liu et al. (2023b); Fu et al. (2019), leaf, text width=16em] ] [Heterogeneous Graph [Liu et al. (2019b); Keerthi Chandra et al. (2020),
Wang et al. (2021d); Zhang et al. (2022a), leaf, text width=16em] ] ] ] [Alignment-based Data
@@
(Sec.4.3.1) [Yin et al. (2020); Zhang et al. (2021b), GSNet Wang et al. (2021a),
ERNIE-GeoL Huang et al. (2022a), Xi et al. (2022), RankETPA Wen et al. (2023b),
Yuan et al. (2021a); Yan et al. (2023), MVMT-STNcite Wang et al. (2023c),
SAInf Ma et al. (2023b), leaf, text width=22em] ] [Encoder-based Alignment
(Sec.4.3.2) [DeepUrbanEvent Jiang et al. (2019), DeepTransport Song et al. (2019b),
Jiang et al. (2019); Cai et al. (2023), RegionDCL Li et al. (2023d), leaf, text width=22em] ] ] [Contrast-based Data
@@
Fusion (Sec.4.4.1) [Li et al. (2022b); Bai et al. (2023); Zhang et al. (2023c),
Mao et al. (2022), leaf, text width=18em] ] [Batch Contrast-based
Fusion (Sec.4.4.2) [Li et al. (2022b); Mao et al. (2022), RecMVC Zhang et al. (2023c),
Bai et al. (2023), UrbanCLIP Yan et al. (2023), leaf, text width=18em] ] ] [Generation-based Data
@@
(Sec.4.5.1) [GeoMAN Liang et al. (2018), Zhao et al. (2021b), leaf, text width=22em ] ] [Mask Modeling-based
Fusion (Sec.4.5.2) [SatMAE Cong et al. (2022), ERNIE-GeoL Huang et al. (2022a), QUERT Xie et al. (2023), G2PTL Wu et al. (2023a),
MGeo Ding et al. (2023), Scale-MAE Reed et al. (2023), leaf, text width=22em ] ] [Diffusion-based Fusion
(Sec.4.5.3) [DiffusionSat Khanna et al. (2023), ChatTraffic Zhang et al. (2023a), DiffUFlow Zheng et al. (2023c),
DiffTraj Yuanshao et al. (2023), leaf, text width=22em ] ] [LLM-enhanced Data
Fusion (Sec.4.5.4) [GeoLLM Manvi et al. (2023), LLM-Mob Wang et al. (2023d), RSGPT Hu et al. (2023a), GeoChat Kuckreja et al. (2023),
LLM4TS Chang et al. (2023), TimeLLM Jin et al. (2023b), UniTime Liu et al. (2023e), leaf, text width=22em ] ] ] ]
@@
DeepST Zhang et al. (2016) Transportation Microsoft 2016
ST-ResNet Zhang et al. (2018) 🍀 Transportation Microsoft 2018
ST-MetaNet+ Pan et al. (2020) Transportation JD Research 2020
DeepCrime Huang et al. (2018) Social JD Research 2021
STUKG Wang et al. (2021d) Transportation THU 2021
DeepSTN+ Lin et al. (2019) Transportation THU 2019
DeepTP Yuan et al. (2021a) Transportation THU 2021
Guo et al. (2019) Transportation BUAA 2019
Photo2Trip Zhao et al. (2017) Transportation USTC/RU 2017
ST-SHN Xia et al. (2021) Public Safety SCUT/HKU 2021
GeoMAN Liang et al. (2018) 🍀 General XDU 2018
Huang et al. (2023a) Urban planning PKU 2023
Liang et al. (2021) Transportation NUS 2021
Balsebre et al. (2022) Urban Planning NTU 2022
Ruan et al. (2022) Transportation NTU 2022
Liu et al. (2023b) Economy HKUST(GZ) 2023
PANDA You et al. (2022) 🍀 Public Safety XMU 2022
UVLens Chen et al. (2021b) Urban Planning XMU 2021
Miyazawa et al. (2019) Transportation SUSTech 2019
NodeSense2Vec Chandra et al. (2021) Social UCF 2021
Keerthi Chandra et al. (2020) Urban Planning UCF 2020
Fu et al. (2019) General UCF 2019
Liu et al. (2022c) Social Gatech 2022
Yuan et al. (2021a) Transportation RMIT 2021
Bai et al. (2019b) 🍀 Transportation Shanghai AI Lab 2019
Ke et al. (2021) Transportation Alibaba 2021
Geng et al. (2019b) Transportation Alibaba 2019
Yao et al. (2018) Transportation PSU 2018
Gao et al. (2022c) Transportation SWJTU 2022
DeepMob Song et al. (2017) Public Safety SUSTech 2017
Feature Based Data Fusion Geng et al. (2019a) Transportation HKUST 2019
Xi et al. (2022) General THU 2022
Zhang et al. (2021b) General THU 2021
Yuan et al. (2021a) 🍀 Transportation THU 2021
Yin et al. (2020) 🍀 Urban Planning NUS 2020
GSNet Wang et al. (2021a) 🍀 Public Safety BJTU 2021
Hashem et al. (2023) General NTU 2023
TrajGAT Yao et al. (2022) Transportation NTU 2022
RADAR Chen et al. (2018) General XMU 2018
Wang et al. (2021e) 🍀 General CSU 2021
Tedjopurnomo et al. (2021) Transportation RMIT 2021
ERNIE-GeoL Huang et al. (2022a) General Baidu 2022
SAInf Ma et al. (2023b) 🍀 Transportation JD Research 2023
Alignment Based Data Fusion Gao et al. (2023a) Transportation SWJTU 2023
 KnowCL Liu et al. (2023g) Economy THU 2023
 Li et al. (2022b) Economy THU 2022
 MMGR Bai et al. (2023) General NTU 2023
 ReMVC Zhang et al. (2023c) Urban Planning NTU 2022
 HMTRL Liu et al. (2022a) Transportation UCF 2023
 Mao et al. (2022) Transportation Shanghai AI Lab 2022
Contrast Baed Data Fusion  UrbanSTC Qu et al. (2022) Transportation JD Research 2022
 UrbanCLIP Yan et al. (2023) General HKUST(GZ) 2023
 SG-GAN Zhang et al. (2020b) Urban Planning NUS 2020
 ActSTD Yuan et al. (2022) Transportation THU 2022
 DiffSTG Wen et al. (2023c) General BJTU 2023
 CP-Route Wen et al. (2023a) Transportation BJTU 2023
 G2PTL Wu et al. (2023a) Transportation Cainiao 2023
 DiffUFlow Zheng et al. (2023c) Transportation CSU 2023
 DP-TFI Xu et al. (2023e) 🍀 Transportation UEST 2023
Wang et al. (2021b) Urban Planning UCF 2021
ChattraffcZhang et al. (2023a) Transportation BJUT 2023
Generation Based Data Fusion  MGEO Ding et al. (2023) General Alibaba 2023
Feature-Based Data Fusion is a straightforward fusion strategy, involving the combination of features from diverse sources such as sensor, visual, and textual data. The core idea is to consolidate raw or processed data features from various sources, forming comprehensive characteristics of studied urban objects. This integration is crucial for enhancing the predictive capabilities of urban models, facilitating complex analyses such as traffic projections and socioeconomic pattern recognition. The distinctive feature of this category lies in directly merging features using methods such as addition, multiplication, concatenation, or graph-based operations.
Alignment-based Data Fusion aims to identify a shared feature space or structure among diverse data representations, enabling their alignment or integration. This involves the model learning to transform information from one source to another for semantic consistency. For instance, in tasks merging images and text, the model must align visual features with textual descriptions, often achieved through a multi-modal embedding space Rasenberg et al. (2020); Baltrušaitis et al. (2018). The widely used attention mechanism Vaswani et al. (2017) exemplifies alignment-based data fusion, assigning weights to different input parts to prioritize them. In tasks like image captioning, cross-modal attention aligns and weights specific image regions with the textual description Luo et al. (2021b); Zhang et al. (2023b); Zohourianshahzadi and Kalita (2022).
Contrast-based Data Fusion employs the contrastive learning framework to enhance feature discriminability at the sample level, contrasting with alignment-based data fusion that focuses on feature-level alignment Deldari et al. (2022); Nakada et al. (2023); Le-Khac et al. (2020). By training the model to differentiate categories or samples, it identifies key distinguishing features. In urban computing, this involves comparing traffic patterns or environmental variables across different areas, time points, and modalities. The pairing of positive and negative samples enables the model to learn and excel in distinguishing complex urban situations, refining the acuity of computational tools Zhang et al. (2023g); Tang et al. (2023); Pan et al. (2023).
Generation-based Data Fusion utilizes deep learning’s creative capacity to generate one urban modality under the condition of other modalities Liu (2023); Fisch et al. (2013); Tran et al. (2021). In urban computing, this approach proves beneficial for simulating diverse scenarios, such as crafting traffic patterns under various circumstances and assessing urban planning outcomes Yuan et al. (2023); Amirian et al. (2019). The generative methods involved in urban multi-modal fusion encompass mask modeling, diffusion, and LLM-enhanced techniques.
@@ 4.2 Feature-based Data Fusion
Feature-based fusion integrates information from different modalities through methods like feature addition, emphasizing equal importance, and feature multiplication, highlighting joint significance (Sec.4.2.1). Additionally, techniques such as feature concatenation (Sec.4.2.2) and graph-based fusion (Sec.4.2.3) enhance multi-modal understanding by combining feature vectors or leveraging graph structures to represent inter-modal connections.
@@ 4.2.1 Feature Addition and Multiplication
For example, Guo et al. (2019) developed the ROD-Revenue framework to predict driver revenue using a linear regression model, given features extracted from multi-source urban data. The basic features from multiple datasets, including ride-on-demand (RoD) service, taxi service, and POI information, are used to construct composite features in a product form. Likewise, Bai et al. (2019b) fused the historical passenger demand, meteorological data, and time meta to learn a joint representation for citywide passenger demand prediction.
@@ 4.2.2 Feature Concatenation
For instance, ST-ResNet Zhang et al. (2017, 2018) dynamically aggregated the outputs of three different residual temporal networks, assigning different weights to different branches and regions. The aggregation was further combined with external factors such as meteorological information (illustrated in Figure 8). Similarly, Lin et al. (2019) proposed a DeepSTN+ framework to forecast urban crowd flows via long-range spatial dependence modeling and the introduction of prior knowledge such as POI distribution. The three categories of historical crowd flow maps - closeness, period and trend, are concatenated firstly, followed by convolution as the fusion of different kinds of information. Some studies Liang et al. (2019b); Ouyang et al. (2020); Sun et al. (2020) followed this concatenation scheme to fuse multimodal information (i.e., external factors) as well. Furthermore, Liang et al. (2021) appended the same three temporal sequences with meta features representing POI and road network (illustrated in Figure 9), which serve as fine-grained knowledge for urban flow prediction.
Figure 8: ST-ResNet dynamically fused the results from three distinct residual temporal networks, allocating varied weights to diverse branches and regions. This fusion process was additionally integrated with external factors, including meteorological information. Zhang et al. (2017, 2018).
In addition, concatenation operation is often used in visual feature fusion. Huang et al. (2023b) concatenated the high-dimensional features of satellite imagery, street-level imagery, and taxi trajectory time series together, and then passed them through a softmax layer to distinguish between urban and non-urban villages. Yin et al. (2023b) also concatenated latent representations of satellite imagery and GPS trace as a fusion method for missing road attribute inference. In addition, the features of household capacity, human mobility, and commercial hotness can be merged together to estimate the population Chen et al. (2021b).
Figure 9: The feature-based data fusion framework proposed by Liang et al. (2021). Features extracted from the temporal sequence are concatenated together with external features from POI and road networks.
@@ 4.2.3 Graph-based Data Fusion
where V represents a set of nodes and E is the set of edges. For a spatio-temporal graph, where the node attributes vary dynamically over time, $\textbf{G}^{(t)}$ is the graph representation at time step $t$ and $\textbf{X}^{(t)}$ is the node feature matrix of graph G at the same time step. Deep learning models based on spatio-temporal graphs have gained significant importance and achieved remarkable success across various applications Yu et al. (2017); Li et al. (2017b); Jain et al. (2016); Wu et al. (2019b).
@@
An intuitive approach for graph-based data fusion is multi-view graph network-based data fusion, which represents diverse data through a set of multi-view graphs Liu et al. (2023b); Fu et al. (2019); Du et al. (2019); Geng et al. (2019a). In this framework, each view is associated with different edge vectors, denoted as $\mathbf{E^{i}}$, which represent the relationships between each node (such POI), denoted as $\mathbf{V}$, under a particular feature view $\mathbf{i}$ (distance, mobility, semantic, etc.). For graphs from different views, graph nodes $\mathbf{V}$ usually serve as connectors, linking them together for feature fusion. Each view in these frameworks captures a unique perspective or aspect of the underlying data, allowing for a comprehensive understanding of a more complex system.
@@
Based on the conception of a multi-view graph, Fu et al. (2019) proposed multi-view POI graph networks that incorporate various geo-features including regions, distances, and human mobility connectivity. Du et al. (2019) developed a group of POI networks to characterize both static and dynamic features through the variable graph edges. Both of these work effectively integrate and fuse information from various data sources and achieve great success in downstream tasks. Liu et al. (2023b) conducted research on characterizing urban vibrancy by employing a multi-view graph framework and demonstrated the effectiveness of multi-view graphs in capturing the intricate relationship between urban vibrancy and urban spatiotemporal dynamics. In addition, different regions in a city can be represented through multi-view graphs as well. Geng et al. (2019a) proposed a multiple graph framework to encode pair-wise correlations between regions in urban areas. As shown in Figure 10, each region square, measuring ($1km\times 1km$), is represented as a node in the graphs. By utilizing graph convolutional networks to observe and fuse these graphs, the framework can capture various region correlations and achieve impressive results in forecasting ride-hailing demand.
@@
Figure 10: The overall architecture of the spatio-temporal multi-graph convolution network proposed by Geng et al. (2019a).
@@
While it is efficient to incorporate features from different data sources using multi-view graphs, describing high-dimensional cross-modal correlation between different items can be challenging as there is no direct interaction between nodes from each graph. To address these difficulties, researchers proposed spatiotemporal heterogeneous networks (SHNs) to encapsulate the complex cross-modal relationships among different urban entities. Subsequently, heterogeneous graph-based data fusion is considered as a promising approach to the fusion of various data in one graph and tackles the above challenges Liu et al. (2019b); Zhang et al. (2022a); Wang et al. (2021d); Keerthi Chandra et al. (2020). Liu et al. (2022c) proposed a hierarchical embedding framework that aimed to establish connections between nodes in the city activity graph and user interaction graph. Although the specific emphasis on SHNs may not be explicitly stressed, this attempt, along with the successful cross-modal node connection, has demonstrated significant success in capturing and representing cross-modal connections. Zhang et al. (2022a) proposed a compound multi-linear relationship graph network to converge various edge connectivity with interaction between multi-view graphs. In their methodology, which is shown in Figure 11, any random walk path in the compound graph is a compound of various kinds of relationships for different views. This work strongly develops the generality of graph neural networks in cross-domain data fusion.
@@
Figure 11: In (a), the multi-view graph connectivity is shown for each individual graph. Vertices represent regions, and the weighted edges represent region-wise relationships. There is no interaction between the graphs. In (b), the compound graph connectivity is represented as a multi-linear graph. Vertices are connected if there is an edge in any of the traditional multi-view graphs. This allows for a more comprehensive understanding of the interdependencies among regions across all graphs Zhang et al. (2022a).
@@
Chandra et al. (2021) designed a SHNs embedding framework that takes the POIs as nodes and human mobility between POIs as weighted links. The SHNs in their research are represented as
@@
In addition to the commonly researched SHNs, Liu et al. (2019b) proposed a multi-modal transportation graph consisting of nodes for users, transport modes, and origin-destination pairs. This approach aimed to develop a multi-modal planning methodology in a transport recommend system. Similarly, Wang et al. (2021d) constructed an urban knowledge graph that effectively integrates features and knowledge from various trajectory data sources to model users’ mobility patterns.

### v2

@@ 4 Methodology Perspective
Multiplication (Sec.4.2.1) [DeepST [363], Bai et al. [12], ROD-Revenue [94], leaf, text width=20em] ] [Feature Concatenation
(Sec.4.2.2) [ST-ResNet [362], DeepCrime [108], DeepSTN+ [170],
ST-MetaNet+ [218], Liang et al. [164], Chen et al. [40],
ST-SHN [311], Huang et al. [112], Yin et al. [337], leaf, text width=20em] ] [Graph-based Feature
Fusion (Sec.4.2.3) [Multi-view Graph [Du et al. [63], Geng et al. [82],
Liu et al. [172], Fu et al. [74], leaf, text width=16em] ] [Heterogeneous Graph [Liu et al. [176], Keerthi Chandra et al. [134],
Wang et al. [278], Zhang et al. [365], leaf, text width=16em] ] ] ] [Alignment-based Data
@@
(Sec.4.3.1) [Yin et al. [339], Zhang et al. [368], GSNet [272],
ERNIE-GeoL [110], Xi et al. [309], RankETPA [298],
Yuan et al. [349], Yan et al. [323], MVMT-STNcite [288],
SAInf [193], leaf, text width=22em] ] [Encoder-based Alignment
(Sec.4.3.2) [DeepUrbanEvent [123], DeepTransport [247],
Jiang et al. [123], Cai et al. [28], RegionDCL [156], leaf, text width=22em] ] ] [Contrast-based Data
@@
Fusion (Sec.4.4.1) [Li et al. [154], Bai et al. [10], Zhang et al. [366],
Mao et al. [196], leaf, text width=18em] ] [Batch Contrast-based
Fusion (Sec.4.4.2) [Li et al. [154], Mao et al. [196], RecMVC [366],
Bai et al. [10], UrbanCLIP [323], leaf, text width=18em] ] ] [Generation-based Data
@@
(Sec.4.5.1) [GeoMAN [162], Zhao et al. [382], leaf, text width=22em ] ] [Mask Modeling-based
Fusion (Sec.4.5.2) [SatMAE [49], ERNIE-GeoL [110], QUERT [314], G2PTL [304],
MGeo [61], Scale-MAE [230], leaf, text width=22em ] ] [Diffusion-based Fusion
(Sec.4.5.3) [DiffusionSat [136], ChatTraffic [358], DiffUFlow [401],
DiffTraj [356], leaf, text width=22em ] ] [LLM-enhanced Data
Fusion (Sec.4.5.4) [GeoLLM [195], LLM-Mob [291], RSGPT [105], GeoChat [141],
LLM4TS [34], Time-LLM [126], UniTime [181], leaf, text width=22em ] ] ] ]
@@
4.1 Overview
The integration of deep learning techniques into urban computing has facilitated the development of various deep-learning-based data fusion methods. These techniques aim to leverage the inherent connections within diverse urban data streams. To better understand the differences in the underlying concepts of data fusion in these studies, we consult prior surveys on methodologies in Urban Computing [390, 177] and taxonomies of deep learning methods in other fields [282, 348, 56, 76] and categorize the lasted researches into four distinct groups based on their underlying techniques, as illustrated in Figure 8. The definitions of each category are outlined as follows:
@@ 4.1 Overview
DeepST [363] Transportation Microsoft 2016
ST-ResNet [364] 🍀 Transportation Microsoft 2018
ST-MetaNet+ [218] Transportation JD Research 2020
DeepCrime [108] Social JD Research 2021
STUKG [278] Transportation THU 2021
DeepSTN+ [170] Transportation THU 2019
DeepTP [349] Transportation THU 2021
Guo et al. [94] Transportation BUAA 2019
Photo2Trip [386] Transportation SU/RU/UCA 2017
ST-SHN [311] Public Safety SCUT/HKU 2021
GeoMAN [162] 🍀 General XDU 2018
Huang et al. [112] Urban planning PKU 2023
Liang et al. [164] Transportation NUS 2021
Balsebre et al. [13] Urban Planning NTU 2022
Ruan et al. [235] Transportation NTU 2022
Liu et al. [172] Economy HKUST(GZ) 2023
PANDA [342] 🍀 Public Safety XMU 2022
UVLens [40] Urban Planning XMU 2021
Miyazawa et al. [200] Transportation SUSTech 2019
NodeSense2Vec [33] Social UCF 2021
Keerthi Chandra et al. [134] Urban Planning UCF 2020
Fu et al. [74] General UCF 2019
Liu et al. [184] Social Gatech 2022
Yuan et al. [349] Transportation RMIT 2021
Bai et al. [12] 🍀 Transportation Shanghai AI Lab 2019
Ke et al. [132] Transportation Alibaba 2021
Geng et al. [84] Transportation Alibaba 2019
Yao et al. [329] Transportation PSU 2018
Gao et al. [79] Transportation SWJTU 2022
DeepMob [249] Public Safety SUSTech 2017
Feature Based Data Fusion Geng et al. [82] Transportation HKUST 2019
Xi et al. [309] General THU 2022
Zhang et al. [368] General THU 2021
Yuan et al. [349] 🍀 Transportation THU 2021
Yin et al. [339] 🍀 Urban Planning NUS 2020
GSNet [272] 🍀 Public Safety BJTU 2021
Hashem et al. [98] General NTU 2023
TrajGAT [328] Transportation NTU 2022
RADAR [39] General XMU 2018
Wang et al. [287] 🍀 General CSU 2021
Tedjopurnomo et al. [261] Transportation RMIT 2021
ERNIE-GeoL [110] General Baidu 2022
SAInf [193] 🍀 Transportation JD Research 2023
Alignment Based Data Fusion Gao et al. [77] Transportation SWJTU 2023
 KnowCL [186] Economy THU 2023
 Li et al. [154] Economy THU 2022
 MMGR [10] General NTU 2023
 ReMVC [366] Urban Planning NTU 2022
 HMTRL [173] Transportation UCF 2023
 Mao et al. [196] Transportation Shanghai AI Lab 2022
Contrast Baed Data Fusion  UrbanSTC [226] Transportation JD Research 2022
 UrbanCLIP [323] General HKUST(GZ) 2023
 SG-GAN [378] Urban Planning NUS 2020
 ActSTD [354] Transportation THU 2022
 DiffSTG [299] General BJTU 2023
 CP-Route [297] Transportation BJTU 2023
 G2PTL [304] Transportation Cainiao 2023
 DiffUFlow [401] Transportation CSU 2023
 DP-TFI [320] 🍀 Transportation UESTC 2023
Wang et al. [273] Urban Planning UCF 2021
Chattraffc[358] Transportation BJUT 2023
Generation Based Data Fusion  MGEO [61] General Alibaba 2023
Feature-Based Data Fusion is a straightforward fusion strategy, involving the combination of features from diverse sources such as sensor, visual, and textual data. The core idea is to consolidate raw or processed data features from various sources, forming comprehensive characteristics of studied urban objects. This integration is crucial for enhancing the predictive capabilities of urban models, facilitating complex analyses such as traffic projections and socioeconomic pattern recognition. The distinctive feature of this category lies in directly merging features using methods such as addition, multiplication, concatenation, or graph-based operations. This type of fusion has relatively low computational complexity, primarily depending on the feature dimensions and the specific fusion operation, typically linear $O(n)$. Here, $n$ represents the number of feature dimensions involved in the fusion process, indicating total number of features being combined from various sources.
Alignment-based Data Fusion aims to identify a shared feature space or structure among diverse data representations, enabling their alignment or integration. This involves the model learning to transform information from one source to another for semantic consistency. For instance, in tasks merging images and text, the model must align visual features with textual descriptions, often achieved through a multi-modal embedding space [229, 14]. The widely used attention mechanism [268] exemplifies alignment-based data fusion, assigning weights to different input parts to prioritize them. In tasks like image captioning, cross-modal attention aligns and weights specific image regions with the textual description [190, 361, 406]. The computational complexity could be higher due to extensive matrix operations, typically quadratic $O(n^{2})$, where $n$ is the input sequence length.
Contrast-based Data Fusion employs the contrastive learning framework to enhance feature discriminability at the sample level, contrasting with alignment-based data fusion that focuses on feature-level alignment [56, 202, 147]. By training the model to differentiate categories or samples, it identifies key distinguishing features. In urban computing, this involves comparing traffic patterns or environmental variables across different areas, time points, and modalities. The pairing of positive and negative samples enables the model to learn and excel in distinguishing complex urban situations, refining the acuity of computational tools [373, 258, 216]. Utilizing contrastive learning to enhance feature discriminability, this method requires significant computation for sample pairing and similarity calculations, with complexity often quadratic $O(n^{2})$. Here, $n$ represents the number of samples involved in the contrastive learning process, indicating the total number of data points being compared.
Generation-based Data Fusion utilizes deep learning’s creative capacity to generate one urban modality under the condition of the same or other modalities [179, 71, 265]. In urban computing, this approach proves beneficial for simulating diverse scenarios, such as crafting traffic patterns under various circumstances and assessing urban planning outcomes [355, 5]. The generative methods involved in urban multi-modal fusion encompass mask modeling, diffusion, and LLM-enhanced techniques. Employing generative models such as GANs and VAEs to simulate urban scenarios, this approach involves highly complex training and optimization processes, leading to very high computational complexity, also generally quadratic $O(n^{2})$ or higher. Specifically, $n$ indicates the number of data points or parameters involved in the training and optimization process of the generative model, indicating the scale of the data used to simulate urban scenarios.
@@ 4.2 Feature-based Data Fusion
Feature-based fusion integrates information from different modalities through methods like feature addition, emphasizing equal importance, and feature multiplication, highlighting joint significance (Sec.4.2.1). Additionally, techniques such as feature concatenation (Sec.4.2.2) and graph-based fusion (Sec.4.2.3) enhance multi-modal understanding by combining feature vectors or leveraging graph structures to represent inter-modal connections. This fusion approach is characterized by its simplicity and efficiency, making it particularly advantageous for applications requiring high real-time performance, such as transportation. However, it is relatively coarse and lacks the ability to capture and align details accurately, which may lead to underperformance in tasks necessitating the amalgamation of diverse datasets, such as urban planning.
@@ 4.2.1 Feature Addition and Multiplication
For example, Guo et al. [94] developed the ROD-Revenue framework to predict driver revenue using a linear regression model, given features extracted from multi-source urban data. The basic features from multiple datasets, including ride-on-demand (RoD) service, taxi service, and POI information, are used to construct composite features in a product form. Likewise, Bai et al. [12] fused the historical passenger demand, meteorological data, and time meta to learn a joint representation for citywide passenger demand prediction.
@@ 4.2.2 Feature Concatenation
For instance, ST-ResNet [362, 364] dynamically aggregated the outputs of three different residual temporal networks, assigning different weights to different branches and regions. The aggregation was further combined with external factors such as meteorological information (illustrated in Figure 9). Similarly, Lin et al. [170] proposed a DeepSTN+ framework to forecast urban crowd flows via long-range spatial dependence modeling and the introduction of prior knowledge such as POI distribution. The three categories of historical crowd flow maps - closeness, period and trend, are concatenated firstly, followed by convolution as the fusion of different kinds of information. Some studies [163, 213, 253] followed this concatenation scheme to fuse multimodal information (i.e., external factors) as well. Furthermore, Liang et al. [164] appended the same three temporal sequences with meta features representing POI and road network (illustrated in Figure 10), which serve as fine-grained knowledge for urban flow prediction.
Figure 9: ST-ResNet dynamically fused the results from three distinct residual temporal networks, allocating varied weights to diverse branches and regions. This fusion process was additionally integrated with external factors, including meteorological information. [362, 364].
In addition, concatenation operation is often used in visual feature fusion. Huang et al. [113] concatenated the high-dimensional features of satellite imagery, street-level imagery, and taxi trajectory time series together, and then passed them through a softmax layer to distinguish between urban and non-urban villages. Yin et al. [337] also concatenated latent representations of satellite imagery and GPS trace as a fusion method for missing road attribute inference. In addition, the features of household capacity, human mobility, and commercial hotness can be merged together to estimate the population [40].
Figure 10: The feature-based data fusion framework proposed by Liang et al. [164]. Features extracted from the temporal sequence are concatenated together with external features from POI and road networks.
@@ 4.2.3 Graph-based Data Fusion
where V represents a set of nodes and E is the set of edges. For a spatio-temporal graph, where the node attributes vary dynamically over time, $\textbf{G}^{(t)}$ is the graph representation at time step $t$ and $\textbf{X}^{(t)}$ is the node feature matrix of graph G at the same time step. Deep learning models based on spatio-temporal graphs have gained significant importance and achieved remarkable success across various applications [343, 158, 115, 308].
@@
An intuitive approach for graph-based data fusion is multi-view graph network-based data fusion, which represents diverse data through a set of multi-view graphs [172, 74, 63, 82]. In this framework, each view is associated with different edge vectors, denoted as $\mathbf{E^{i}}$, which represent the relationships between each node (such POI), denoted as $\mathbf{V}$, under a particular feature view $\mathbf{i}$ (distance, mobility, semantic, etc.). For graphs from different views, graph nodes $\mathbf{V}$ usually serve as connectors, linking them together for feature fusion. Each view in these frameworks captures a unique perspective or aspect of the underlying data, allowing for a comprehensive understanding of a more complex system.
@@
Based on the conception of a multi-view graph, Fu et al. [74] proposed multi-view POI graph networks that incorporate various geo-features including regions, distances, and human mobility connectivity. Du et al. [63] developed a group of POI networks to characterize both static and dynamic features through the variable graph edges. Both of these work effectively integrate and fuse information from various data sources and achieve great success in downstream tasks. Liu et al. [172] conducted research on characterizing urban vibrancy by employing a multi-view graph framework and demonstrated the effectiveness of multi-view graphs in capturing the intricate relationship between urban vibrancy and urban spatiotemporal dynamics. In addition, different regions in a city can be represented through multi-view graphs as well. Geng et al. [82] proposed a multiple graph framework to encode pair-wise correlations between regions in urban areas. As shown in Figure 11, each region square, measuring ($1km\times 1km$), is represented as a node in the graphs. By utilizing graph convolutional networks to observe and fuse these graphs, the framework can capture various region correlations and achieve impressive results in forecasting ride-hailing demand.
@@
Figure 11: The overall architecture of the spatio-temporal multi-graph convolution network proposed by Geng et al. [82].
@@
While it is efficient to incorporate features from different data sources using multi-view graphs, describing high-dimensional cross-modal correlation between different items can be challenging as there is no direct interaction between nodes from each graph. To address these difficulties, researchers proposed spatiotemporal heterogeneous networks (SHNs) to encapsulate the complex cross-modal relationships among different urban entities. Subsequently, heterogeneous graph-based data fusion is considered as a promising approach to the fusion of various data in one graph and tackles the above challenges [176, 365, 278, 134, 408, 292]. Liu et al. [184] proposed a hierarchical embedding framework that aimed to establish connections between nodes in the city activity graph and user interaction graph. Although the specific emphasis on SHNs may not be explicitly stressed, this attempt, along with the successful cross-modal node connection, has demonstrated significant success in capturing and representing cross-modal connections. Zhang et al. [365] proposed a compound multi-linear relationship graph network to converge various edge connectivity with interaction between multi-view graphs. In their methodology, which is shown in Figure 12, any random walk path in the compound graph is a compound of various kinds of relationships for different views. This work strongly develops the generality of graph neural networks in cross-domain data fusion.
@@
Figure 12: In (a), the multi-view graph connectivity is shown for each individual graph. Vertices represent regions, and the weighted edges represent region-wise relationships. There is no interaction between the graphs. In (b), the compound graph connectivity is represented as a multi-linear graph. Vertices are connected if there is an edge in any of the traditional multi-view graphs. This allows for a more comprehensive understanding of the interdependencies among regions across all graphs [365].
@@
Chandra et al. [33] designed a SHNs embedding framework that takes the POIs as nodes and human mobility between POIs as weighted links. The SHNs in their research are represented as
@@
In addition to the commonly researched SHNs, Liu et al. [176] proposed a multi-modal transportation graph consisting of nodes for users, transport modes, and origin-destination pairs. This approach aimed to develop a multi-modal planning methodology in a transport recommend system. Similarly, Wang et al. [278] constructed an urban knowledge graph that effectively integrates features and knowledge from various trajectory data sources to model users’ mobility patterns.